# OccScene: Semantic Occupancy-based Cross-task Mutual Learning for 3D Scene Generation

### v1

@@ II-A Diffusion Models for Scene Generation
Diffusion Models, a recently established class of generative models grounded in non-equilibrium thermodynamics theory [30], which delineate empirical data distributions through an iterative noise reduction mechanism [31], closely paralleling score-based generative models that rely on Langevin dynamics leveraging inferred data distribution gradients [8]. Diffusion models have significantly advanced fields such as text-to-image generation [7, 32, 33, 34] and controllable video generation [35, 36, 37, 38]. These models have also been developed to support downstream applications, notably in autonomous driving scene generation [12, 20, 22, 39, 11]. One class of the driving generators is based on NeRF and Gaussian Splatting [40, 41, 42, 43], which suffer from poor diversity. Another class involves world models or world generators, with notable examples including DriveDreamer [12], BEVGen [44], Panacea [45], Drive-WM [22], etc. Recently, some studies have utilized generated data to enhance downstream perception models. DetDiffusion [15] introduces perception-aware loss and attributes to improve the quality of the generation images for 2D object detection. To facilitate realistic scene generation, MagicDrive [21] leverages 3D geometry information from ground-truth labels. However, this method depends heavily on ground-truth labels in the inference process and faces significant challenges in generating flexible and generalizable real-world scenes.
@@ II-B Semantic Occupancy Prediction
Semantic Occupancy Prediction (SOP) is a dense 3D perception task that unifies semantic segmentation with scene completion [46, 47, 5]. Prior research has extensively employed LiDAR to capitalize on its 3D geometric data capabilities [48, 1, 49]. Recently, camera-driven 3D SOP has garnered significant interest due to the affordability and mobility of camera systems [50, 51, 52, 53, 23, 54, 55]. MonoScene [23] propose to infer both geometry and semantics from a single RGB image via 2D-to-3D feature projection, which sparked a wave of advancements in camera-based scene understanding [56, 57, 58, 5]. TPVFormer [56] innovates with a tri-perspective framework for detailed 3D scene depiction. OccFormer [57] devises a dual-path transformer to handle dense 3D feature processing for semantic occupancy. SurroundOcc [58] introduces multi-view image inputs for enhanced occupancy estimation. Pioneered by VPD [9], conditional diffusion models are leveraged for 3D perception tasks including multi-view stereo and semantic occupancy prediction. However, how to use the powerful generative models to produce high-quality data pairs and thus improve perception remains unexplored.
@@
Figure 2: Overview of the proposed OccScene. The framework involves the concurrent training of the perception model and the generative diffusion UNet, facilitating the simultaneous generation of images or videos and their corresponding semantic occupancy grids during the inference process. Within OccScene, we introduce a Mamba-based Dual Alignment (MDA) module to sequentially align the semantic occupancy and the diffusion latent with camera trajectory awareness.
@@
7:    $\quad{\nabla_{\theta}\left\|f_{\theta}\left(\bm{X}_{(occ,text)},\sqrt{\bar{\alpha}_{t}}\bm{y}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\bar{\alpha}_{t}\right)-\epsilon\right\|}+{\nabla_{\delta}\left\|\bm{X}_{occ}-\bm{\tilde{X}}_{occ}\right\|}$
@@ III-A Preliminaries
The standard generative diffusion models aim to establish one-to-many mappings with a forward and reverse process [59]. In the forward process, the input image $\bm{y}_{0}$ is progressively corrupted to $\bm{y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in ${T}$ time steps following a discrete-time Markov chain. The distributions of intermediate steps can be characterized by marginalizing: $q\left(\bm{y}_{t}\mid\bm{y}_{0}\right)=\mathcal{N}\left(\bm{y}_{t}\mid\sqrt{\bar{\alpha}_{t}}\bm{y}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}$ is a pre-defined coefficient. $\mathcal{N}$ and $\mathbf{I}$ denote the normal distribution and the identity matrix, respectively. In the reverse process, a diffusion neural network such as UNet estimates a corresponding image to approximate the input image $\bm{y}_{0}$ from noisy input $\bm{y}_{T}$ and conditions $\bm{X}$ which are optionally provided to guide the estimation process. Each step of the reverse process can be defined as conditional distribution transition [59], which is formulated as: $p_{\theta}\left(\bm{y}_{0:T}\mid\bm{X}\right)=p\left(\bm{y}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\bm{y}_{t-1}\mid\bm{y}_{t},\bm{X}\right)$, where $p_{\theta}$ represents the reverse function and $\bm{X}$ denotes the conditions of the diffusion model.
@@ III-D Analysis on the Benefits of Cross-task Mutual Learning
Leveraging these insights, we observe that cross-task mutual learning facilitates the discovery of higher-quality solutions characterized by more robust minima. As depicted in Figure 5, we conducted extensive experiments to assess the impact of mutual learning. Our findings indicate that the framework performs better on training data when employing cross-task mutual learning, resulting in a more stable learning process and a better minimum of the training loss, as shown in Figure 5(a). Specifically, while both independent and mutual learning initially reduce loss values, the independent learning strategy tends to stagnate in local minima during the middle stages of training, hindering its ability to converge effectively. In contrast, the mutual learning curve exhibits a steady decline throughout training, suggesting the identification of a broader and more optimal minimum, which is indicative of improved performance [67, 68]. Figure 5(b) further illustrates that through the joint learning scheme, our framework enables cross-task collaboration, leading to enhanced performance in both generation and perception tasks.
@@ IV-B Datasets and Evaluation Metrics
NYUv2. The NYUv2 dataset [50] comprises 1449 indoor scenes captured via Kinect, represented as 240$\times$144$\times$240 occupancy grids annotated with 13 distinct classes: 11 semantic categories, alongside labels for free space and unknown areas. The accompanying RGBD input has a resolution of 640×480 pixels. Following [23], we employ 795 instances for training and 654 instances for testing.
SemanticKITTI. The SemanticKITTI dataset [46] includes 22 diverse outdoor scenes featuring both LiDAR scans and stereo image pairs. Its ground truth is structured into 256$\times$256$\times$32 occupancy grids, each measuring 0.2m in all dimensions and annotated with 21 semantic categories, including 19 specific semantics, one class for free space, and another for unknown areas.
@@
ControlNet [12] 448 $\times$ 640 $50.61$ 192 $\times$ 512 $65.24$
SD (Finetune) [7] 448 $\times$ 640 $47.82$ 192 $\times$ 512 $60.55$
@@
MonoScene[23]$+$ours 10.77 93.48 15.72 17.74 15.76 48.44 40.33 18.45 17.33 32.39 17.14 29.78
@@
MonoScene[23]$+$ours 62.59 33.20 22.65 3.41 19.40 26.67 14.27 1.85 2.07 7.00 22.54 5.11 39.95 4.42 1.46 0.00 8.08 6.42 3.49 14.98
TPVFormer [56]$+$ours 61.93 32.95 20.89 0.21 23.24 32.46 15.97 2.41 1.98 8.82 26.20 8.04 33.27 2.52 0.87 0.00 9.39 10.90 6.80 15.73
TABLE IV: Quantitative results on the NuScenes-Occupancy validation set. The $**{}^{\text{*}}$ and $**{}^{\text{*}}$ take images and LiDAR-projected depth maps as inputs. The top two performers are marked bold and underline.
@@
$**{}^{\text{*}}$ [75] 11.5 4.0 11.8 12.3 5.1 3.8 6.2 6.0 8.2 7.5 24.1 13.0 12.8 11.5 11.6 20.2 10.6
$**{}^{\text{*}}$ [76] 12.0 5.1 10.7 12.4 6.5 4.0 5.0 6.3 8.0 7.2 21.8 14.8 13.0 11.8 12.0 21.2 10.7
MonoScene[23]$+$ours 9.7 6.4 11.0 12.6 6.3 7.1 8.3 4.4 7.3 10.7 24.8 15.1 17.1 14.7 8.8 11.8 11.0
TPVFormer [56]$+$ours 10.4 7.3 12.4 13.5 9.6 10.9 10.7 5.0 7.6 11.8 24.9 17.5 16.8 15.2 9.3 12.6 12.2
@@ IV-C Main Results
Generation Evaluation. The quantitative results of scene generation are presented in Figure 6. For the baseline models of SD [7], we fine-tune it on the datasets with the same training setting as OccScene. ControlNet [12] is conditioned with semantic maps and depth maps jointly to generate corresponding images. As illustrated in Figure 6, both SD and ControlNet tend to generate unreasonable geometry (e.g., cars in columns 3 and 4) and blurred details (e.g., distant structures in columns 1 and 2), especially in complex scenes and distant regions. We also present the quantitative results of cross-view generation in Figure 7. The model of tune-a-video [35] is further fine-tuned on the SemanticKITTI dataset. Compared to existing models, our method generates more consistent and reasonable results across different perspectives. The significant superiority stems from the incorporated fine-grained geometry and semantics as the perception priors.
Moreover, we report the quantitative results with different datasets in Table I(a) and Table I(b). Our proposed OccScene outperforms other methods in terms of image and video generation with equals or exceeds baseline resolution of 256 $\times$ 448, achieving 113.28 FVD on SemanticKITT [46] and 11.87 FID on NuScenes-Occupancy [70]. The extensive experiments demonstrate the effectiveness of our proposed method on high-fidelity scene generation.
@@
Learning Process Analysis. The results generated from different denoising steps are visualized in Figure 8. As the generated images become clearer, the semantic occupancy predicted by the perception model becomes more complete and accurate. As discussed in Section III-B, to ensure the stable training of the perception model, the varying scales of $\sqrt{\bar{\alpha}_{t}}$ are employed in the loss supervision. As shown in Figure 9, such operation significantly improves the stability of the loss curve and enhances the perception performance with higher mIoU scores.
@@
TABLE V: Comparison about support for Semantic Occupancy Prediction model (i.e., MonoScene). Results are reported by testing on the NuScenes-Occupancy validation set..
Training Support for Semantic Occupancy Prediction. As shown in Table V, we conduct a training support experiment to demonstrate that OccScene can generate synthetic image-occupancy data pairs to enhance the training for the perception task of Semantic Occupancy Prediction. To produce the data pairs, we generate the same amount of images as the original dataset. Note that the Semantic Occupancy Prediction model of MonoScene [23] is trained from scratch on the synthetic data to enable fair comparisons. As shown in the table, OccScene significantly improves MonoScene in terms of both IoU and mIoU for Semantic Occupancy Prediction. While the compared method of MagicDrive [21] only marginally improves mIoU. We attribute such a difference to the high fidelity and fine-grained geometric control of generated results from OccScene.
@@
w/ JDS&\&&MDA 19.86 113.28 15.70
TABLE VI: Ablation studies of the framework components on the SemanticKITTI validation set. The ‘JDS’ and ‘MDA’ denote the Joint Diffusion Scheme and the Mamba-based Dual Alignment, respectively.

### v2

@@ II-A Diffusion Models for Scene Generation
Diffusion Models, a recently established class of generative models grounded in non-equilibrium thermodynamics theory [30], which delineate empirical data distributions through an iterative noise reduction mechanism [31], closely paralleling score-based generative models that rely on Langevin dynamics leveraging inferred data distribution gradients [8]. Diffusion models have significantly advanced fields such as text-to-image generation [7, 32, 33, 34] and controllable video generation [35, 36, 37, 38]. These models have also been developed to support downstream applications, notably in autonomous driving scene generation [12, 20, 22, 39, 11]. One class of the driving generators is based on NeRF and Gaussian Splatting [40, 41, 42, 43], which suffer from poor diversity. Another class involves world models or world generators, with notable examples including DriveDreamer [12], BEVGen [44], Panacea [45], Drive-WM [22], etc. Recent advancements in LiDAR generation and semantic scene generation also highlighted the potential of diffusion models. In LiDiff [46], a diffusion model is adapted to directly process sparse 3D LiDAR point clouds for scene completion, achieving superior detail recovery compared to range-image-based methods. SemCity [47] introduces a triplane diffusion model for semantic scene generation to address data sparsity challenges in real-world outdoor environments. Recently, some studies have utilized generated data to enhance downstream perception models. DetDiffusion [15] introduces perception-aware loss and attributes to improve the quality of the generation images for 2D object detection. To facilitate realistic scene generation, MagicDrive [21] leverages 3D geometry information from ground-truth labels. However, this method depends heavily on ground-truth labels in the inference process and faces significant challenges in generating flexible and generalizable real-world scenes.
@@ II-B Semantic Occupancy Prediction
Semantic occupancy prediction(SOP) is a dense 3D perception task that unifies semantic segmentation with scene completion [48]. Prior research has extensively employed LiDAR to capitalize on its 3D geometric data capabilities [49, 50, 51, 52, 53]. SSCNet [49] pioneers an end-to-end 3D convolutional network for joint occupancy and semantic prediction from a single depth image. LMSCNet [50] introduces lightweight multiscale architectures for efficient scene completion. JS3CNet [51] leverages contextual shape priors from sequential LiDAR data to enhance sparse point cloud segmentation. SCPNet [52] improves robustness through innovative sub-network designs and knowledge distillation. PaSCo [53] extends semantic scene completion to panoptic scene completion, introducing uncertainty awareness critical for robotics applications. Recent self-supervised methods have also advanced occupancy prediction [54, 55, 56]. SceneRF [54] employs neural radiance fields (NeRF) with explicit depth optimization for monocular 3D reconstruction, excelling in novel depth synthesis. Behind the Scenes [55] proposes a density field-based approach for volumetric occupancy prediction, effectively handling occlusions. SelfOcc [56] introduces a self-supervised framework using video sequences, eliminating the need for voxel annotations by leveraging signed distance fields (SDF) and multi-view constraints. Moreover, camera-driven 3D SOP has garnered significant interest due to the affordability and mobility of camera systems [57, 51, 58, 59, 23, 60, 61]. MonoScene [23] propose to infer both geometry and semantics from a single RGB image via 2D-to-3D feature projection, which sparked a wave of advancements in camera-based scene understanding [62, 63, 64, 5]. TPVFormer [62] innovates with a tri-perspective framework for detailed 3D scene depiction. OccFormer [63] devises a dual-path transformer to handle dense 3D feature processing for semantic occupancy. SurroundOcc [64] introduces multi-view image inputs for enhanced occupancy estimation. Pioneered by VPD [9], conditional diffusion models are leveraged for 3D perception tasks including multi-view stereo and semantic occupancy prediction. However, how to use the powerful generative models to produce high-quality data pairs and thus improve perception remains unexplored.
@@
Figure 2: Overview of the proposed OccScene. The framework involves the concurrent training of the perception model and the generative diffusion UNet. A reconstruction loss is leveraged between the ground-truth images or videos $\bm{y}$ and generation results $\bm{y}_{0}$, while a perception loss is adopted between the ground-truth semantic occupancy $\bm{\tilde{X}}_{occ}$ and predicted semantic occupancy $\bm{X}_{occ}$. During the inference process, OccScene takes Gaussian noise $\bm{y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and conditional text prompt $\bm{X}_{text}$ as inputs, facilitating the simultaneous generation of images or videos $\bm{y}_{0}$ and their associated semantic occupancy $\bm{X}_{occ}$. Within OccScene, we introduce a Mamba-based Dual Alignment (MDA) module to sequentially align the semantic occupancy and the diffusion latent with camera trajectory awareness.
@@
7:  $\quad{\nabla_{\theta}\left\|f_{\theta}\left(\bm{X}_{(occ,text)},\sqrt{\bar{\alpha}_{t}}\bm{y}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\bar{\alpha}_{t}\right)-\epsilon\right\|}+{\nabla_{\delta}\left\|\bm{\tilde{X}}_{occ}-\bm{X}_{occ}\right\|}$
@@ III-A Preliminaries
The standard generative diffusion models aim to establish one-to-many mappings with a forward and reverse process [65]. In the forward process, the input image $\bm{y}_{0}$ is progressively corrupted to $\bm{y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in T{T} time steps following a discrete-time Markov chain. The distributions of intermediate steps can be characterized by marginalizing: $q\left(\bm{y}_{t}\mid\bm{y}_{0}\right)=\mathcal{N}\left(\bm{y}_{t}\mid\sqrt{\bar{\alpha}_{t}}\bm{y}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}$ is a pre-defined coefficient. $\mathcal{N}$ and $\mathbf{I}$ denote the normal distribution and the identity matrix, respectively. In the reverse process, a diffusion neural network such as UNet estimates a corresponding image to approximate the input image $\bm{y}_{0}$ from noisy input $\bm{y}_{T}$ and conditions $\bm{X}$ which are optionally provided to guide the estimation process. Each step of the reverse process can be defined as conditional distribution transition [65], which is formulated as: $p_{\theta}\left(\bm{y}_{0:T}\mid\bm{X}\right)=p\left(\bm{y}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\bm{y}_{t-1}\mid\bm{y}_{t},\bm{X}\right)$, where $p_{\theta}$ represents the reverse function and $\bm{X}$ denotes the conditions of the diffusion model.
@@ III-D Analysis on the Benefits of Cross-task Mutual Learning
Leveraging these insights, we observe that cross-task mutual learning facilitates the discovery of higher-quality solutions characterized by more robust minima. As depicted in Figure 5, we conducted extensive experiments to evaluate the impact of mutual learning. Our findings indicate that the framework performs better on training data when cross-task mutual learning is employed, leading to a more stable learning process and a better minimum of the training loss, as shown in Figure 5(a). Specifically, while both independent and mutual learning initially reduce the loss, the independent learning strategy tends to stagnate in local minima during the middle stages of training, limiting its ability to converge effectively. In contrast, the mutual learning curve exhibits a steady decline throughout training, suggesting the identification of a broader and more optimal minimum, which indicates improved performance [73, 74]. Figure 5(b) further illustrates that through the joint learning scheme, our framework enables cross-task collaboration, leading to enhanced performance in both generation and perception tasks.
@@ IV-B Datasets and Evaluation Metrics
[…]
SemanticKITTI. The SemanticKITTI dataset [75] includes 22 diverse outdoor scenes featuring both LiDAR scans and stereo image pairs. Its ground truth is structured into $\times256$$\times$32 occupancy grids, each measuring 0.2m in all dimensions and annotated with 21 semantic categories, including 19 specific semantics, one class for free space, and another for unknown areas.
@@
ControlNet [12] 448 $\times$ 640 $50.61$ 192 $\times$ 512 $65.24$ -
SD (Finetune) [7] 448 $\times$ 640 $47.82$ 192 $\times$ 512 $60.55$ -
@@
NDC-Scene [83] 12.02 93.51 13.11 13.77 15.83 49.57 39.87 17.17 24.57 31.00 14.96 29.03
ISO [84] 14.21 93.47 15.89 15.14 18.35 50.01 40.82 18.25 25.90 34.08 17.67 31.25
MonoScene[23]++ours 10.77 93.48 15.72 17.74 15.76 48.44 40.33 18.45 17.33 32.39 17.14 29.78
NDC-Scene [83]++ours 15.06 94.85 16.51 16.97 17.84 51.10 42.41 19.56 26.33 34.51 18.14 32.12
@@
OccFormer [63] 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86 13.46
CGFormer [85] 65.51 32.31 20.82 0.16 23.52 34.32 19.44 4.61 2.71 7.67 26.93 8.83 39.54 2.38 4.08 0.00 9.20 10.67 7.84 16.87
MonoScene[23]++ours 62.59 33.20 22.65 3.41 19.40 26.67 14.27 1.85 2.07 7.00 22.54 5.11 39.95 4.42 1.46 0.00 8.08 6.42 3.49 14.98
TPVFormer [62]++ours 61.93 32.95 20.89 0.21 23.24 32.46 15.97 2.41 1.98 8.82 26.20 8.04 33.27 2.52 0.87 0.00 9.39 10.90 6.80 15.73
OccFormer [63]++ours 65.73 33.96 20.22 1.20 23.65 34.85 26.58 3.37 2.79 10.63 27.85 8.97 36.86 3.00 4.22 0.00 10.36 11.25 5.74 17.43
[…]
@@
[…]
MonoScene[23]++ours 9.7 6.4 11.0 12.6 6.3 7.1 8.3 4.4 7.3 10.7 24.8 15.1 17.1 14.7 8.8 11.8 11.0
TPVFormer [62]++ours 10.4 7.3 12.4 13.5 9.6 10.9 10.7 5.0 7.6 11.8 24.9 17.5 16.8 15.2 9.3 12.6 12.2
[…]
@@ IV-C Main Results
Generation Evaluation. The quantitative results of scene generation are presented in Figure 6. For the baseline models of SD [7], we fine-tune it on the datasets with the same training setting as OccScene. ControlNet [12] is conditioned with semantic maps and depth maps jointly to generate corresponding images. As illustrated in Figure 6, both SD and ControlNet tend to generate unreasonable geometry (e.g., cars in columns 3 and 4) and blurred details (e.g., distant structures in columns 1 and 2), especially in complex scenes and distant regions. We also present the quantitative results of cross-view generation in Figure 7. The model of tune-a-video [35] is further fine-tuned on the SemanticKITTI dataset. Compared to existing models, our method generates more consistent and reasonable results across different perspectives. The significant superiority stems from the incorporated fine-grained geometry and semantics as the perception priors. Moreover, we report the quantitative results with different datasets in Table I(a) and Table I(b). Our proposed OccScene outperforms other methods in terms of image and video generation with equals or exceeds baseline resolution of 256 $\times$ 448, achieving 113.28 FVD on SemanticKITT [75] and 11.87 FID on NuScenes-Occupancy [77]. The extensive experiments demonstrate the effectiveness of our proposed method on high-fidelity scene generation.
[…]
SSD [86] 112.82 0.12 2.23
SemCity [47] 56.55 0.04 3.25
OccScene (Ours) 39.21 0.02 4.17
TABLE V: Quantitative comparison of 3D semantic scene generation performance on the SemanticKITTI validation set. The top two performers are marked bold and underline. Our proposed method outperforms previous works in terms of the 3D semantic scene generation quality.
Figure 8: Quantitative comparison of 3D semantic scene generation with SemCity [47]. Our method efficiently generates more complete scenes with detailed structures, especially in road surfaces and vehicle shapes.
The qualitative evaluation results of 3D semantic scene generation are shown in Table V. Following SemCity [47], we assess generation quality using Frechet Inception Distance (FID) [78], Kernel Inception Distance (KID) [87], and Inception Score (IS). Our proposed OccScene outperforms previous works in terms of generation quality, achieving a 30.66% improvement in FID compared to SemCity [47] and a 65.25% improvement compared to SSD [86]. These results demonstrate the effectiveness of our approach for high-fidelity 3D scene generation. The quantitative evaluation results of 3D semantic scene generation are illustrated in Figure 8. Note that we independently replicated the semantic scene generation model of SemCity [47] following their official implementations, as no pretrained weights are available in the public repository. Facilitated by the cross-task mutual benefits, our OccScene produces more realistic and complete 3D scene generation results, especially in overall completion (e.g., road surfaces in the first three columns) and structural details (e.g., vehicle shapes in the fourth column).
@@
Learning Process Analysis. The results generated from different denoising steps are visualized in Figure 9. As the generated images become clearer, the semantic occupancy predicted by the perception model becomes more complete and accurate. As discussed in Section III.B, to stabilize training, the loss function incorporates $\sqrt{\bar{\alpha}_{t}}$ to adaptively scale supervision signals according to noise levels. As shown in Figure 10, this modification significantly improves the stability of the loss curve and enhances the perception performance, as evidenced by higher mIoU scores.
@@
TABLE VI: Comparison about support for Semantic Occupancy Prediction model (i.e., MonoScene). Results are reported by testing on the NuScenes-Occupancy validation set..
Training Support for Semantic Occupancy Prediction. As shown in Table VI, we conduct a training support experiment to demonstrate that OccScene can generate synthetic image-occupancy data pairs to enhance the training for the perception task of Semantic Occupancy Prediction. To produce the data pairs, we generate the same amount of images as the original dataset. Note that the Semantic Occupancy Prediction model of MonoScene [23] is trained from scratch on the synthetic data to enable fair comparisons. As shown in the table, OccScene significantly improves MonoScene in terms of both IoU and mIoU for Semantic Occupancy Prediction. While the compared method of MagicDrive [21] only marginally improves mIoU. We attribute such a difference to the high fidelity and fine-grained geometric control of generated results from OccScene.
@@
w/ JDS&\&MDA 19.86 113.28 14.98
TABLE VII: Ablation studies of the framework components on the SemanticKITTI validation set. The ‘JDS’ and ‘MDA’ denote the Joint Diffusion Scheme and the Mamba-based Dual Alignment, respectively.